WHAT IS THE ANSWER OF (2X+5) (3X-2) = (6X+4) (X-8)-33
Answer by rothauserc(4718)

(2X+5) (3X-2) = (6X+4) (X-8)-33
6X^2 +11X -10 = 6X^2 -44X -32 -33
divide both sides of = by 6X^2
11X -10 = -44X -65
add 44X to both sides of =
55X -10 = -65
add 10 to both sides of =
55X = -55
X = -1
check answer by substituting for X in original equation
(2*(-1)+5) (3(-1)-2) = (6(-1)+4) (-1-8) - 33
3 * -5 = -2 * -9 -33
-15 = 18 -33
-15 = -15
our answer checks out
X = -1